# Haoyuan Li,  Mathias Funk, Aaqib Saeed

### v1

@@ 1 Introduction
Federated Learning (FL) holds immense promise for privacy-centric collaborative AI, yet its practical deployment is complex (Luping et al., 2019; Wu et al., 2020; Liu et al., 2021a). Designing an effective FL system requires navigating a range of challenges, including statistical heterogeneity, systems constraints, and shifting task objectives (Marfoq et al., 2021; Jhunjhunwala et al., 2022; Lu et al., 2024). To date, this design process has been a manual, labor-intensive effort led by domain experts, resulting in static, bespoke solutions that are brittle in the face of real-world dynamics. This paper argues that this manual design paradigm is the critical bottleneck hindering the widespread adoption of FL (Li et al., 2020b; Linardos et al., 2022; Tang et al., 2023). We propose to address this bottleneck by introducing Helmsman, a new frontier focused on the automated synthesis of FL systems using autonomous AI agents. While recent breakthroughs have produced LLM-based agents capable of remarkable feats in general-purpose code generation (Zhang et al., 2024; Tao et al., 2024a; Islam et al., 2024; Nunez et al., 2024; Tao et al., 2024b; Novikov et al., 2025), their ability to reason about and systematically construct the complex, multi-component systems required for robust federated learning remains a fundamental open challenge.
@@
The difficulty of FL design is rooted in its vast and combinatorial nature. Real-world deployments rarely present a single, isolated challenge. Instead, they involve a confluence of issues: clients may have non-IID data (Li et al., 2020a; Reddi et al., 2020), heterogeneous computational capabilities (Diao et al., 2020; Kim et al., 2023), and unreliable network connections (Chen et al., 2020; Liu et al., 2024a), all while pursuing diverse task objectives (Marfoq et al., 2021; Lu et al., 2024). Current FL research often operates in silos, developing point solutions for individual problems. While effective in isolation, these solutions are difficult to compose, and their interactions are unpredictable. This leaves practitioners facing an intractable design space, underscoring the need for an automated approach that can reason holistically about multifaceted constraints and synthesize integrated, deployment-ready solutions.
@@
While the need for automation is clear, existing agent-based code generation frameworks are ill-equipped for the unique demands of FL. Systems like those proposed by Muennighoff et al. (2023) and Dong et al. (2024), often built on single-agent architectures with prompting techniques like Chain-of-Thought (Wei et al., 2022) or ReAct (Yao et al., 2023), excel at solving self-contained programming tasks. However, they struggle when confronted with the system-level complexity of FL. The challenge is not merely to generate a correct algorithm, but to orchestrate an entire distributed system comprising interdependent modules for data handling, client-side training, server-side aggregation, and overall strategy. This requires reasoning about algorithmic trade-offs, resource constraints, and communication protocols simultaneously—a holistic design task that exceeds the capacity of monolithic, single-agent approaches and is amplified by the need for a realistic simulation environment for validation.
@@ 2.1 Related Work
A key insight from recent agentic AI research is that complex problem-solving is often best addressed not by a single monolithic agent, but by a collaborative ensemble of specialized agents. This multi-agent paradigm, rooted in the principle of “division of labor,” has proven highly effective in automated code generation. For instance, systems like AgentCoder (Huang et al., 2023) separate the roles of code implementation and test case generation, while CodeSim (Islam et al., 2025) introduces a simulation-driven process to verify plans and debug code internally. This approach of decomposing a complex task—such as emulating the human programming cycle of retrieval, planning, coding, and debugging (Islam et al., 2024)—consistently enhances the robustness and quality of the generated solution.
The power of this collaborative approach extends beyond self-contained coding tasks to system-level engineering and scientific discovery. Multi-agent teams have demonstrated superior performance in resolving real-world GitHub issues in benchmarks like SWE-bench (Jimenez et al., 2024; Chen et al., 2024; Liu et al., 2024b), navigating complex engineering design spaces (Ni & Buehler, 2024; Ocker et al., 2025), and even automating scientific discovery by generating and testing novel hypotheses (Gottweis et al., 2025; Naumov et al., 2025). These successes highlight a clear trend: as the complexity of a task grows, so too do the benefits of a structured, multi-agent approach. However, despite its proven potential, this paradigm has not yet been systematically applied to the unique, multifaceted challenges of designing and deploying Federated Learning systems. This is the critical gap our work addresses.
@@ 2.2 Motivation
Combinatorial Complexity of Strategies. Real-world FL problems are rarely one-dimensional. As illustrated in Figure 2, a typical deployment requires addressing a combination of challenges spanning data heterogeneity, system constraints, robustness, and more. For each challenge, a plethora of specialized algorithms exist—FedProx for stragglers (Li et al., 2020a), SCAFFOLD for client drift (Karimireddy et al., 2020), FedNova for system heterogeneity (Wang et al., 2020), etc. Manually selecting, combining, and tuning these strategies for a specific problem is a combinatorial task that quickly becomes intractable for a human expert, highlighting the need for a system that can automate this exploration.
Brittleness in Dynamic Environments. Hand-crafted FL solutions are often static and optimized for a specific set of assumptions about client data, network conditions, and model architectures. Consequently, they are brittle and behave unpredictably when deployed into dynamic, real-world environments where these conditions may fluctuate (Wu et al., 2022; Li et al., 2021). For instance, a strategy may be sensitive to the number of participating clients (Jhunjhunwala et al., 2022; Xu et al., 2021) or the degree of resource heterogeneity (Diao et al., 2020; Kim et al., 2023). An automated system must be able to design solutions that are not only performant but also robust to such environmental variations.
Dichotomy in FL Frameworks. The FL ecosystem is split between two poles. On one hand, research frameworks like Flower (Beutel et al., 2020) and PySyft (Ziller et al., 2021) offer the modularity and flexibility essential for rapid prototyping and innovation. On the other, industrial platforms like FATE (Liu et al., 2021b) and NVIDIA FLARE (Roth et al., 2022) prioritize the scalability, robustness, and operational efficiency required for production. This dichotomy creates a gap between experimental research and deployable solutions. An ideal automated system must bridge this gap, combining the innovative flexibility of research frameworks with a process that ensures production-grade reliability.
@@ 3.4 Agent Instrumentation and Tooling
To ensure research plans are both state-of-the-art and technically sound, the Planning Agent leverages a dual-source knowledge system. A Web Search Tool (via Tavily API) provides access to the most current, real-world information, such as up-to-date library documentation and best practices, grounding the plan in practical realities. Complementing this, a Retrieval-Augmented Generation (RAG) pipeline queries a vector database of seminal FL literature from arXiv to retrieve state-of-the-art techniques. This RAG process employs a multi-stage approach—initiating with a hybrid BM25 (Robertson et al., 2009) and vector search for broad recall, followed by a Cohere rerank-v3.5 model (Cohere, 2024) with Voyage-3-large embeddings (VoyageAI, 2025) to ensure high precision—allowing the agent to identify the most relevant algorithms for the user’s specific problem.
@@
The autonomous evaluation and refinement loop (Section 3.3) is critically dependent on a reliable execution environment. For this, Helmsman utilizes the Flower framework (Beutel et al., 2020) as a sandboxed Simulation Tool. This environment is the core of the diagnose-and-repair cycle. It provides the essential ground-truth feedback by executing the generated code and capturing the resulting logs, enabling the Evaluator and Debugger Agents to analyze system behavior and perform targeted corrections.
@@ 4.1 Benchmark Design
Evaluating generative agents like Helmsman  demands a paradigm shift from traditional software and ML benchmarks. Code generation benchmarks such as HumanEval (Chen et al., 2021) assess an agent’s ability to solve self-contained, algorithmic problems. Standard FL benchmarks, conversely, evaluate a pre-defined model’s performance on a static dataset. Neither is sufficient for measuring an agent’s capacity for end-to-end research automation—the ability to take a high-level scientific goal and autonomously synthesize a complete, functional, and performant system.
Table 2: Performance evaluation on heterogeneous federated learning benchmarks (a). We compare our agent-synthesized strategy against task-specific baselines. The best performing method is marked in bold, while the second-best is underlined. Results for specialized methods are denoted by symbols: FedNova (Wang et al., 2020), FedNS (Li et al., 2024), and HeteroFL (Diao et al., 2020).
@@
Q1 Object Recognition CIFAR-10-LT Quantity Skew 70.68{70.68} 69.13{69.13} $\textbf{78.39}^{*}$ $\underline{76.26}$
Q2 Object Recognition CIFAR-100-C Feature Skew 35.29{35.29} 36.93{36.93} $\textbf{42.13}^{\dagger}$ $\underline{39.43}$
Q3 Object Recognition CIFAR-10N Label Noise 74.32{74.32} ${79.42}^{*}$ $\underline{80.67}^{\dagger}$ 81.14
@@
Q4 Object Recognition Office-Home Domain Shift 52.94{52.94} 50.81{50.81} $\textbf{57.68}^{*}$ $\underline{53.47}$
Q5 Human Activity HAR User Heterogeneity 94.38{94.38} 95.66{95.66} $\underline{95.77}^{*}$ 96.74
Q6 Speech Recognition Speech Commands Speaker Variation $\underline{84.93}$ 84.11{84.11} ${83.08}^{*}$ 86.19
Q7 Medical Diagnosis Fed-ISIC2019 Site Heterogeneity 55.43{55.43} 61.79{61.79} $\underline{63.62}^{*}$ 63.80
Q8 Object Recognition Caltech101 Class Imbalance 48.21{48.21} 47.75{47.75} $\underline{63.62}^{*}$ 50.92
@@
Q9 Object Recognition CIFAR-100 Resource Constraint 60.83{60.83} 59.86{59.86} $\underline{62.19}^{\ddagger}$ 63.49
@@
To address this gap, we introduce AgentFL-Bench, a benchmark designed to rigorously evaluate the capabilities of agentic systems in federated learning. To ensure a thorough test of agent versatility, it comprises $16$ unique tasks curated to span five pivotal FL research domains: data heterogeneity, communication efficiency, personalization, active learning, and continual learning (Figure 3). Furthermore, these tasks are designed for realism, reflecting the multifaceted nature of real-world FL challenges where issues like non-IID data and system constraints often co-occur (see Figure 2). This requires the agent to reason about and integrate appropriate strategies, rather than solving isolated toy problems. Finally, to enable consistent and reproducible evaluation, every task is defined by a standardized natural language query (Table 1). This provides an unambiguous problem specification that minimizes prompt-engineering variability and allows for fair, apples-to-apples comparisons between different agentic systems. The complete set of task queries, along with their specific experimental configurations and established baseline implementations, are detailed in Appendix A.1 and A.2.
Table 3: Performance evaluation on heterogeneous federated learning benchmarks (b). We compare our agent-synthesized strategy against task-specific baselines. Results for specialized methods are denoted by symbols: FedNova (Wang et al., 2020), FedPer§ (Arivazhagan et al., 2019), and FedWelt (Yoon et al., 2021).
@@
Q10 Object Recognition CIFAR-100 Bandwidth Limits 42.14{42.14} 45.26{45.26} $\underline{46.17}^{*}$ 48.43
Q11 Handwriting Recog. FEMNIST Connectivity Limits 87.06{87.06} 88.19{88.19} $\underline{88.86}^{*}$ 89.03
@@
Q12 Handwriting Recog. FEMNIST Local Adaptation 74.61{74.61} 75.32{75.32} $\textbf{77.06}^{\S}$ $\underline{76.75}$
Q13 Object Recognition CIFAR-10 Distribution Skew 75.83{75.83} $\underline{80.09}$ $\textbf{81.74}^{\S}$ 79.11{79.11}
@@ 4.2 Implementation Details
Our agentic system, Helmsman, is constructed upon the LangGraph (LangChainAI, 2025) framework, leveraging LangChain (LangChainAI, 2024) for tool integration. The system’s LLM backbone consists of Google’s Gemini-2.5-flash (Comanici et al., 2025) for the planning stage and Claude-Sonnet-4 (Anthropic, 2025) for the coding and evaluation stages. The maximum debugging attempts during the evaluation stage is set to 10 times. We evaluate Helmsman on our AgentFL-Bench  benchmark, which covers cross-silo and cross-device scenarios (Tables 6, 7, and 8). To ensure consistency, all tasks run for 100 communication rounds, with each client performing 5 local updates per round, except where noted in Appendix A.2. Further details on datasets, models, and task-specific configurations are provided in Tables 9 and 10.
To evaluate the efficacy of our system’s strategy module, we perform a comparative analysis against several FL baselines. This is achieved by substituting our generated strategy with a baseline method while maintaining identical task configurations. Our comparison includes foundational baselines, such as FedAvg (McMahan et al., 2017) and FedProx (Li et al., 2020a), alongside specialized methods designed for specific challenges: FedNova (Wang et al., 2020) for distribution shifts, FedNS (Li et al., 2024) for noisy data, HeteroFL (Diao et al., 2020) for heterogeneous models, FedPer (Arivazhagan et al., 2019) for personalization, FAST Li et al. (2025) for active learning, and FedWeIT (Yoon et al., 2021) for continual learning.
Table 4: Performance evaluation on heterogeneous federated learning benchmarks (c). We compare our agent-synthesized strategy against task-specific baselines. Results for specific federated learning methods are denoted by symbols: FedWelt (Yoon et al., 2021), FAST# (Li et al., 2025).
@@
Q14 Medical Diagnosis DermaMNIST Sample Selection 71.70{71.70} 71.51{71.51} $\textbf{74.37}^{\#}$ $\underline{72.64}$
Q15 Object Recognition CIFAR-10 Distribution Skew 64.19{64.19} 66.03{66.03} $\textbf{77.14}^{\#}$ $\underline{68.59}$
@@
Q16 Object Recognition Split-CIFAR100 Incremental Tasks 16.38{16.38} 16.71{16.71} $\underline{28.56}^{\P}$ 51.04

### v2

@@ 1 Introduction
Federated Learning (FL) holds immense promise for privacy-centric collaborative AI, yet its practical deployment is complex (luping2019cmfl; wu2020safa; liu2021fedcpf). Designing an effective FL system requires navigating a range of challenges, including statistical heterogeneity, systems constraints, and shifting task objectives (marfoq2021federated; fedvarp; lu2024fedhca2). To date, this design process has been a manual, labor-intensive effort led by domain experts, resulting in static, bespoke solutions that are brittle in the face of real-world dynamics. This paper argues that this manual design paradigm is the critical bottleneck hindering the widespread adoption of FL (li2020multi; linardos2022federated; tang2023privacy). We propose to address this bottleneck by introducing Helmsman, a new frontier focused on the automated synthesis of FL systems using autonomous AI agents. While recent breakthroughs have produced LLM-based agents capable of remarkable feats in general-purpose code generation (zhang2024pybench; tao2024codelutra; islam2024mapcoder; nunez2024autosafecoder; tao2024magis; novikov2025alphaevolve), their ability to reason about and systematically construct the complex, multi-component systems required for robust federated learning remains a fundamental open challenge.
@@
The difficulty of FL design is rooted in its vast and combinatorial nature. Real-world deployments rarely present a single, isolated challenge. Instead, they involve a confluence of issues: clients may have non-IID data (fedprox; fedopt), heterogeneous computational capabilities (heterofl; depthfl), and unreliable network connections (chen2020asynchronous; liu2024fedasmu), all while pursuing diverse task objectives (marfoq2021federated; lu2024fedhca2). Current FL research often operates in silos, developing point solutions for individual problems. While effective in isolation, these solutions are difficult to compose, and their interactions are unpredictable. This leaves practitioners facing an intractable design space, underscoring the need for an automated approach that can reason holistically about multifaceted constraints and synthesize integrated, deployment-ready solutions.
@@
While the need for automation is clear, existing agent-based code generation frameworks are ill-equipped for the unique demands of FL. Systems like those proposed by muennighoff2023octopack and dong2024self, often built on single-agent architectures with prompting techniques like Chain-of-Thought (wei2022chain) or ReAct (yao2023react), excel at solving self-contained programming tasks. However, they struggle when confronted with the system-level complexity of FL. The challenge is not merely to generate a correct algorithm, but to orchestrate an entire distributed system comprising interdependent modules for data handling, client-side training, server-side aggregation, and overall strategy. This requires reasoning about algorithmic trade-offs, resource constraints, and communication protocols simultaneously—a holistic design task that exceeds the capacity of monolithic, single-agent approaches and is amplified by the need for a realistic simulation environment for validation.
@@ 2.1 Related Work
A key insight from recent agentic AI research is that complex problem-solving is often best addressed not by a single monolithic agent, but by a collaborative ensemble of specialized agents. This multi-agent paradigm, rooted in the principle of “division of labor,” has proven highly effective in automated code generation. For instance, systems like AgentCoder (huang2023agentcoder) separate the roles of code implementation and test case generation, while CodeSim (islam2025codesim) introduces a simulation-driven process to verify plans and debug code internally. This approach of decomposing a complex task—such as emulating the human programming cycle of retrieval, planning, coding, and debugging (islam2024mapcoder)—consistently enhances the robustness and quality of the generated solution.
The power of this collaborative approach extends beyond self-contained coding tasks to system-level engineering and scientific discovery. Multi-agent teams have demonstrated superior performance in resolving real-world GitHub issues in benchmarks like SWE-bench (jimenez2024swebench; chen2024coder; liu2024marscode), navigating complex engineering design spaces (ni2024mechagents; ocker2025idea), and even automating scientific discovery by generating and testing novel hypotheses (gottweis2025towards; naumov2025dora). These successes highlight a clear trend: as the complexity of a task grows, so too do the benefits of a structured, multi-agent approach. However, despite its proven potential, this paradigm has not yet been systematically applied to the unique, multifaceted challenges of designing and deploying Federated Learning systems. This is the critical gap our work addresses.
@@ 2.2 Motivation
Combinatorial Complexity of Strategies. Real-world FL problems are rarely one-dimensional. As illustrated in Figure 2, a typical deployment requires addressing a combination of challenges spanning data heterogeneity, system constraints, robustness, and more. For each challenge, a plethora of specialized algorithms exist—FedProx for stragglers (fedprox), SCAFFOLD for client drift (scaffold), FedNova for system heterogeneity (fednova), etc. Manually selecting, combining, and tuning these strategies for a specific problem is a combinatorial task that quickly becomes intractable for a human expert, highlighting the need for a system that can automate this exploration.
Brittleness in Dynamic Environments. Hand-crafted FL solutions are often static and optimized for a specific set of assumptions about client data, network conditions, and model architectures. Consequently, they are brittle and behave unpredictably when deployed into dynamic, real-world environments where these conditions may fluctuate (fedkd; moon). For instance, a strategy may be sensitive to the number of participating clients (fedvarp; fedcm) or the degree of resource heterogeneity (heterofl; depthfl). An automated system must be able to design solutions that are not only performant but also robust to such environmental variations.
Dichotomy in FL Frameworks. The FL ecosystem is split between two poles. On one hand, research frameworks like Flower (flower) and PySyft (pysyft) offer the modularity and flexibility essential for rapid prototyping and innovation. On the other, industrial platforms like FATE (fate) and NVIDIA FLARE (nvidiaflare) prioritize the scalability, robustness, and operational efficiency required for production. This dichotomy creates a gap between experimental research and deployable solutions. An ideal automated system must bridge this gap, combining the innovative flexibility of research frameworks with a process that ensures production-grade reliability.
@@ 3.4 Agent Instrumentation and Tooling
To ensure research plans are both state-of-the-art and technically sound, the Planning Agent leverages a dual-source knowledge system. A Web Search Tool (via Tavily API) provides access to the most current, real-world information, such as up-to-date library documentation and best practices, grounding the plan in practical realities. Complementing this, a Retrieval-Augmented Generation (RAG) pipeline queries a vector database of seminal FL literature from arXiv to retrieve state-of-the-art techniques. This RAG process employs a multi-stage approach—initiating with a hybrid BM25 (bm25) and vector search for broad recall, followed by a Cohere rerank-v3.5 model (cohere-rerank-3.5) with Voyage-3-large embeddings (voyage-3-large) to ensure high precision—allowing the agent to identify the most relevant algorithms for the user’s specific problem.
@@
The autonomous evaluation and refinement loop (Section 3.3) is critically dependent on a reliable execution environment. For this, Helmsman utilizes the Flower framework (flower) as a sandboxed Simulation Tool. This environment is the core of the diagnose-and-repair cycle. It provides the essential ground-truth feedback by executing the generated code and capturing the resulting logs, enabling the Evaluator and Debugger Agents to analyze system behavior and perform targeted corrections.
@@ 4.1 Benchmark Design
Evaluating generative agents like Helmsman  demands a paradigm shift from traditional software and ML benchmarks. Code generation benchmarks such as HumanEval (humaneval) assess an agent’s ability to solve self-contained, algorithmic problems. Standard FL benchmarks, conversely, evaluate a pre-defined model’s performance on a static dataset. Neither is sufficient for measuring an agent’s capacity for end-to-end research automation—the ability to take a high-level scientific goal and autonomously synthesize a complete, functional, and performant system.
Table 2: Performance evaluation on heterogeneous federated learning benchmarks (a). We compare our agent-synthesized strategy against task-specific baselines. Results are averaged over 3 independent runs with standard deviation. The best performing method is marked in bold, while the second-best is underlined. Results for specialized methods are denoted by symbols: FedNova (fednova), FedNS (fedns), and HeteroFL (heterofl).
@@
Q1 Object Recognition CIFAR-10-LT Quantity Skew $70.08_{\pm 0.52}$} $69.65_{\pm 0.42}$} $\textbf{78.96}_{\pm 0.41}^{*}$ $\underline{76.71}_{\pm 0.43}$
Q2 Object Recognition CIFAR-100-C Feature Skew $33.96_{\pm 1.68}$} $35.13_{\pm 1.80}$} $\textbf{42.79}_{\pm 0.67}^{\dagger}$ $\underline{39.75}_{\pm 0.30}$
Q3 Object Recognition CIFAR-10N Label Noise $73.95_{\pm 1.31}$} $78.78_{\pm 0.49}$} $\underline{80.55}_{\pm 0.47}^{\dagger}$ $\textbf{81.62}_{\pm 0.62}$
@@
Q4 Object Recognition Office-Home Domain Shift $53.02_{\pm 1.03}$} $50.67_{\pm 0.39}$} $\textbf{57.26}_{\pm 0.60}^{*}$ $\underline{54.49}_{\pm 0.97}$
Q5 Human Activity HAR User Heterogeneity $94.84_{\pm 0.44}$} $95.22_{\pm 0.32}$} $\underline{95.19}_{\pm 0.75}^{*}$ $\textbf{96.28}_{\pm 0.42}$
Q6 Speech Recognition Speech Commands Speaker Variation $\underline{84.44}_{\pm 0.36}$ $84.19_{\pm 0.29}$} $83.48_{\pm 0.49}$$^{*}$ $\textbf{86.58}_{\pm 0.38}$
Q7 Medical Diagnosis Fed-ISIC2019 Site Heterogeneity $57.09_{\pm 1.44}$} $61.11_{\pm 0.71}$} $\underline{62.88}_{\pm 0.53}^{*}$ $\textbf{63.75}_{\pm 0.85}$
Q8 Object Recognition Caltech101 Class Imbalance $47.99_{\pm 0.67}$} $47.20_{\pm 0.64}$} $\textbf{63.77}_{\pm 0.27}^{*}$ $\underline{50.88}_{\pm 1.22}$
@@
Q9 Object Recognition CIFAR-100 Resource Constraint $59.96_{\pm 0.93}$} $59.43_{\pm 0.68}$} $\underline{62.62}_{\pm 0.42}^{\ddagger}$ $\textbf{62.94}_{\pm 0.53}$
@@
To address this gap, we introduce AgentFL-Bench, a benchmark designed to rigorously evaluate the capabilities of agentic systems in federated learning. To ensure a thorough test of agent versatility, it comprises $16$ unique tasks curated to span five pivotal FL research domains: data heterogeneity, communication efficiency, personalization, active learning, and continual learning (Figure 3). Furthermore, these tasks are designed for realism, reflecting the multifaceted nature of real-world FL challenges where issues like non-IID data and system constraints often co-occur (see Figure 2). This requires the agent to reason about and integrate appropriate strategies, rather than solving isolated toy problems. Finally, to enable consistent and reproducible evaluation, every task is defined by a standardized natural language query (Table 1). This provides an unambiguous problem specification that minimizes prompt-engineering variability and allows for fair, apples-to-apples comparisons between different agentic systems. The complete set of task queries, along with their specific experimental configurations and established baseline implementations, are detailed in Appendix A.4 and A.5.
Table 3: Performance evaluation on heterogeneous federated learning benchmarks (b). We compare our agent-synthesized strategy against task-specific baselines. Results are averaged over 3 independent runs with standard deviation. Results for specialized methods are denoted by symbols: FedNova (fednova), FedPer§ (fedper), and FedWelt (fedweit).
@@
Q10 Object Recognition CIFAR-100 Bandwidth Limits $41.77_{\pm 0.79}$} $45.21_{\pm 0.61}$} $\underline{45.77}_{\pm 0.39}^{*}$ $\textbf{48.78}_{\pm 0.29}$
Q11 Handwriting Recog. FEMNIST Connectivity Limits $87.46_{\pm 0.43}$} $87.95_{\pm 0.37}$} $\underline{89.11}_{\pm 0.20}^{*}$ $\textbf{89.73}_{\pm 0.73}$
@@
Q12 Handwriting Recog. FEMNIST Local Adaptation $74.12_{\pm 0.38}$} $75.47_{\pm 0.36}$} $\textbf{77.43}_{\pm 0.27}^{\S}$ $\underline{76.41}_{\pm 0.36}$
Q13 Object Recognition CIFAR-10 Distribution Skew $72.59_{\pm 2.61}$} $\underline{79.76}_{\pm 0.27}$ $\textbf{81.56}_{\pm 0.51}^{\S}$ $79.22_{\pm 0.35}$}
@@ 4.2 Implementation Details
Our agentic system, Helmsman, is constructed upon the LangGraph (langgraph) framework, leveraging LangChain (langchain) for tool integration. The system’s LLM backbone consists of Google’s Gemini-2.5-flash (gemini-2.5) for the planning stage and Claude-Sonnet-4.0 (claude-4) for the coding and evaluation stages. We also conduct additional experiments with different LLM choices in the Appendix (Claude-Sonnet-4.5 and GPT-5.1). The maximum debugging attempts during the evaluation stage is set to 10 times. We evaluate Helmsman on our AgentFL-Bench  benchmark, which covers cross-silo (i.e., 5 clients) and cross-device (i.e., 10 clients) scenarios. We report the extensive details of datasets, models, and task-specific configurations for each research query in Tables 16 and 17.
To ensure consistency, all queries run for 100 communication rounds, with each client performing 5 local updates per round. We adopt the 80/20 data split across all FL tasks, where the training data is partitioned and distributed to each client based on the specific research task requirement. To evaluate the efficacy of our system’s strategy module, we perform a comparative analysis against several FL baselines. This is achieved by substituting our generated strategy with a baseline method while maintaining identical task configurations. Our comparison includes foundational baselines, such as FedAvg (fedavg) and FedProx (fedprox), alongside specialized methods designed for specific challenges: FedNova (fednova) for distribution shifts, FedNS (fedns) for noisy data, HeteroFL (heterofl) for heterogeneous models, FedPer (fedper) for personalization, FAST fast for active learning, and FedWeIT (fedweit) for continual learning.
Table 4: Performance evaluation on heterogeneous federated learning benchmarks (c). We compare our agent-synthesized strategy against task-specific baselines. Results are averaged over 3 independent runs with standard deviation. Results for specific federated learning methods are denoted by symbols: FedWelt (fedweit), FAST# (fast).
@@
Q14 Medical Diagnosis DermaMNIST Sample Selection $71.13_{\pm 0.66}$} $71.43_{\pm 0.25}$} $\textbf{74.71}_{\pm 0.41}^{\#}$ $\underline{72.87}_{\pm 0.23}$
Q15 Object Recognition CIFAR-10 Distribution Skew $63.75_{\pm 1.00}$} $66.02_{\pm 0.23}$} $\textbf{77.95}_{\pm 0.66}^{\#}$ $\underline{68.80}_{\pm 0.21}$
@@
Q16 Object Recognition Split-CIFAR100 Incremental Tasks $15.38_{\pm 0.97}$} $15.86_{\pm 0.65}$} $\underline{29.45}_{\pm 0.72}^{\P}$ $\textbf{50.95}_{\pm 1.09}$